# Hubbard and Heisenberg models on hyperbolic lattices - Metal-insulator transitions, global antiferromagnetism and enhanced boundary fluctuations

### v1

@@ Abstract
We study the Hubbard and Heisenberg models on hyperbolic lattices with open boundary conditions by means of mean-field approximations, spin-wave theory and quantum Monte Carlo (QMC) simulations. For the Hubbard model we use the auxiliary-field approach and for Heisenberg systems the stochastic series expansion algorithm. We concentrate on bipartite lattices where the QMC simulations are free of the negative sign problem. The considered lattices are characterized by a Dirac like density of states, Schläfli indices $\{p,q\}=\{10,3\}$ and $\{8,3\}$, as well as by flat bands, $\{8,8\}$. The Dirac density of states cuts off the logarithmic divergence of the staggered spin susceptibility and allows for a finite $U$ semi-metal to insulator transition. This transition has the same mean-field exponents as for the Euclidean counterpart. In the presence of flat bands we observe the onset of magnetic ordering at any finite $U$. The magnetic state at intermediate coupling can be described as a global antiferromagnet. It breaks the $C_{p}$ rotational and time reversal symmetries but remains invariant under combined $C_{p}\mathcal{T}$ transformations. The state is characterized by macroscopic ferromagnetic moments, that globally cancel. We observe that fluctuations on the boundary of the system are greatly enhanced: while spin wave calculations predict the breakdown of antiferromagnetism on the boundary but not in the bulk, QMC simulations show a marked reduction of the staggered moment on the edge of the system.
@@ I Introduction
The effect of dimensionality and point group symmetries on correlation induced phenomena is an established research domain in flat Euclidean space. The aim of this paper is to investigate the effect curvature has on the physics of standard models of correlated electrons and spins. The role of the curvature is widely studied in models of statistical mechanics, like the Ising model [1, 2, 3, 4, 5, 6, 7, 8]. A constant negative curvature can for example lead to the presence of symmetry-broken phases and phase transitions, not present in the flat Euclidean plane [1, 9, 2, 10, 11] or change the critical properties of phase transitions [12, 3, 5]. Also in condensed matter physics the popularity of hyperbolic lattices is increasing. The topics of interest are both single-particle physics, like the electronic band structure [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27] – also under the influence of a magnetic field and topological phenomena – and correlation effects [28, 29, 26]. The motivation can partially be attributed to recent experimental realizations of hyperbolic lattices in circuit quantum electrodynamics and electric circuits [30, 31].
In this paper we study the Hubbard and Heisenberg models on three exemplary regular tessellations of the hyperbolic plane, that exhibit different electronic properties in the non-interacting limit. We use a Hartree-Fock approximation, a spin wave analysis, auxiliary-field and stochastic series expansion quantum Monte Carlo (QMC) methods to investigate the models as a function of the interaction strength and system size. Since all considered lattices are bipartite, we can perform QMC simulations without a negative sign problem for both the Hubbard and Heisenberg models.
The electronic properties of the three lattices resemble a Dirac system with a linear density of states (DOS) around the Fermi level ($\{10,3\}$ and $\{8,3\}$ lattice), and a flat-band system with a high DOS at the Fermi level ($\{8,8\}$ lattice). The mean-field approximation of the Hubbard model on the $\{10,3\}$ lattice is already studied in Ref. [26], and we complement these results with both spin wave calculations and first-time fermionic QMC studies on the hyperbolic lattice. We will compare the spatial distribution of the magnetic ordering in the weak-coupling and strong-coupling limits, the transition from a semi-metal to insulator with antiferromagnetic (AFM) ordering as a function of the interaction strength and thermodynamic properties.
The remainder of the paper is organized as follows. In Sec. II we define the Hubbard and Heisenberg models and the hyperbolic lattices on which we study them. In Sec. III we introduce the mean-field approximation, spin-wave analysis and the QMC algorithms that will be used in this work. In Sec. IV we present our numerical results before we conclude in Sec. V. In the Appendix additional numerical data can be found.
@@ II.1 Hyperbolic lattices
In this work, we consider two-dimensional graphs that can be categorized by their Schläfli index $\{p,q\}$. The lattice sites are located at the corners of regular $p$-gons and every lattice site has coordination number $q$ [32]. Hyperbolic lattices with a constant negative curvature fulfill the condition $(p-2)(q-2)>4$. In contrast, lattices with $(p-2)(q-2)=4$ lie within the flat Euclidean plane and lattices with $(p-2)(q-2)<4$ in a positively curved plane. For the Euclidean plane exactly three solutions exist – the square lattice $\{4,4\}$, the honeycomb lattice $\{6,3\}$ and the triangular lattice $\{3,6\}$, whereas for the hyperbolic case infinitely many pairs $\{p,q\}$ exist that fulfill the condition [33].
@@
We concentrate on finite size hyperbolic lattices with open boundary conditions. The lattice size can be increased by starting with a single $p$-gon and successively adding shells or layers of $p$-gons. The number of lattice sites grows exponentially with the number of layers, strongly limiting the accessible lattice sizes for our calculations. In Fig. 1 the $\{8,3\}$ lattice is shown with up to $n_{l}=3$ layers. In order to properly depict the lattices, the Poincaré disk model is used. All lattice sites are projected into the open unit disk and all edges, connecting two nearest-neighboring sites, are geodesic lines. Those geodesics are circular arcs in the Poincaré disk model, that would hit the unit circle perpendicular and they all have the same length within the hyperbolic metric  [^1] The metric in the hyperbolic plane with a constant negative curvature of -1 is given by $ds^{2}=4\frac{dx^{2}+dy^{2}}{(1-x^{2}-y^{2})^{2}}$. Here, $x$ and $y$ are the coordinates within the open unit circle in the Poincaré disk model. [35, 36].

### v2

# Hubbard and Heisenberg models on hyperbolic lattices: Metal-insulator transitions, global antiferromagnetism, and enhanced boundary fluctuations

@@ Abstract
We study the Hubbard and Heisenberg models on hyperbolic lattices with open boundary conditions by means of mean-field approximations, spin-wave theory, and quantum Monte Carlo (QMC) simulations. For the Hubbard model we use the auxiliary-field approach and for Heisenberg systems the stochastic series expansion algorithm and concentrate on bipartite lattices where the QMC simulations are free of the negative sign problem. The hyperbolic lattices have an extensive number of sites on the boundary, such that one has to distinguish between bulk and total density of states. The considered lattices are characterized by a Dirac-like total density of states, Schläfli indices $\{p,q\}=\{10,3\}$ and $\{8,3\}$, as well as by flat bands, $\{8,8\}$. The Dirac total density of states cuts off the logarithmic divergence of the staggered spin susceptibility and allows for a finite $U$ metal-to-insulator transition. This transition has the same mean-field exponents as for the Gross-Neveu transition in Euclidean space. We argue that this transition is induced by the open-boundary conditions and that it will be absent in the periodic case. In the presence of flat bands we observe the onset of magnetic ordering at any finite $U$. This conclusion holds even though the bulk density of states is constant at the Fermi energy for the three considered lattices. The magnetic state at intermediate coupling can be described as a global antiferromagnet. It breaks the $C_{p}$ rotational and time-reversal symmetries but remains invariant under combined $C_{p}\mathcal{T}$ transformations. The state is characterized by macroscopic ferromagnetic moments, that globally cancel. We observe that fluctuations on the boundary of the system are greatly enhanced: While spin-wave calculations predict the breakdown of antiferromagnetism on the boundary but not in the bulk, QMC simulations show a marked reduction of the staggered moment on the edge of the system.
@@ I Introduction
The effect of dimensionality and point group symmetries on correlation induced phenomena is an established research domain in flat Euclidean space. The aim of this paper is to investigate the effect curvature has on the physics of standard models of correlated electrons and spins. The role of the curvature is widely studied in models of statistical mechanics, like the Ising model [1, 2, 3, 4, 5, 6, 7, 8]. A constant negative curvature can for example lead to the presence of symmetry-broken phases and phase transitions, not present in the flat Euclidean plane [1, 9, 2, 10, 11] or change the critical properties of phase transitions [12, 3, 5]. Also in condensed matter physics the popularity of hyperbolic lattices is increasing. The topics of interest are both single-particle physics, like the electronic band structure [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]—also under the influence of a magnetic field and topological phenomena—and correlation effects [28, 29, 26]. The motivation can partially be attributed to recent experimental realizations of hyperbolic lattices in circuit quantum electrodynamics and electric circuits [30, 31].
In this paper we study the Hubbard and Heisenberg models on three exemplary regular tessellations of the hyperbolic plane, that exhibit different electronic properties in the noninteracting limit. We use a Hartree-Fock approximation, a spin-wave analysis, auxiliary-field and stochastic series expansion quantum Monte Carlo (QMC) methods to investigate the models as a function of the interaction strength and system size. Since all considered lattices are bipartite, we can perform QMC simulations without a negative sign problem for both the Hubbard and Heisenberg models. There is a plethora of exotic phases of matter, such as spin liquids, that have been realized on two-dimensional bipartite lattices [32, 33, 34, 35]. This provides a further motivation to study correlation effects on bipartite hyperbolic lattices. We will concentrate on systems with open boundary conditions. This is important since the number of boundary sites is extensive, such that bulk and total density of states (DOS) differ. The mean-field approximation of the Hubbard model on the $\{10,3\}$ lattice is already studied in Ref. [26], and we complement these results with both spin-wave calculations and first-time fermionic QMC studies on the hyperbolic lattice.
The remainder of the paper is organized as follows. In Sec. II we define the Hubbard and Heisenberg models and the hyperbolic lattices on which we study them. In Sec. III we introduce the mean-field approximation, spin-wave analysis and the QMC algorithms that will be used in this work. In Sec. IV we present our numerical results before we conclude in Sec. V. Additional numerical data as well as details and derivations of the methods used in this paper can be found in the Appendix.
@@ II.1 Hyperbolic lattices
In this work, we consider two-dimensional graphs that can be categorized by their Schläfli index $\{p,q\}$. The lattice sites are located at the corners of regular $p$-gons and every lattice site has coordination number $q$ [36]. Hyperbolic lattices with a constant negative curvature fulfill the condition $(p-2)(q-2)>4$. In contrast, lattices with $(p-2)(q-2)=4$ lie within the flat Euclidean plane and lattices with $(p-2)(q-2)<4$ on a sphere, that has a constant positive curvature. For the Euclidean plane exactly three solutions exist—the square lattice $\{4,4\}$, the honeycomb lattice $\{6,3\}$ and the triangular lattice $\{3,6\}$, and for the sphere five—the five Platonic solids, whereas for the hyperbolic case infinitely many pairs $\{p,q\}$ exist that fulfill the condition [37].
@@
We concentrate on finite-size hyperbolic lattices with open boundary conditions. The lattice size can be increased by starting with a single $p$-gon and successively adding shells or layers of $p$-gons. The number of lattice sites grows exponentially with the number of layers, strongly limiting the accessible lattice sizes for our calculations. In Fig. 1 the $\{8,3\}$ lattice is shown with up to $n_{l}=3$ layers. To properly depict the lattices, the Poincaré disk model is used. All lattice sites are projected into the open unit disk and all edges, connecting two nearest-neighboring sites, are geodesic lines. Those geodesics are circular arcs in the Poincaré disk model, that would hit the unit circle perpendicularly and they all have the same length within the hyperbolic metric  [^1] The metric in the hyperbolic plane with a constant negative curvature of $-1$ is given by $ds^{2}=4\frac{dx^{2}+dy^{2}}{(1-x^{2}-y^{2})^{2}}$. Here, $x$ and $y$ are the coordinates within the open unit circle in the Poincaré disk model. [39, 40].